### v2

@@ 3.2 Assumptions and Main Identification Result
There are at least $d_{X}$ linearly independent points in the support of $X_{it}$ a.s. given $V_{it}$ and $W_{i}$.
Assumption 4 guarantees that $\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]$ in (3.6) is invertible [^12] This is true because for any random vector $X\in\mathbb{R}^{d_{X}}$, $\mathbb{E}\left[XX^{{}^{\prime}}\right]$ is singular if and only if there exists a non-random vector $a\in\mathbb{R}^{d_{X}}\backslash\left\{0\right\}$ such that $a^{{}^{\prime}}X=0$ almost surely., thus the APE and LAR can be identified by (3.6) and the LIE. As discussed by Imbens and Newey (2009), requiring residual variation in $X_{it}$ conditioning on $V_{it}$ is generally not restrictive. Essentially, it requires $Z_{it}$ to move $X_{it}$ sufficiently well such that the set $\left\{X_{it}:V_{it}=v\text{ for }v\in\text{Supp}\left(V_{it}\right)\right\}$ is not a singleton. Therefore, it is similar to the relevance condition in the classical IV literature. [^13] I thank an anonymous referee for pointing this out. The restriction on the support of $X_{it}$ mainly comes from conditioning on $W_{i}$. When parametric justifications for Assumption 2 are used (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2024)) so that $W_{i}$ only includes averages of $X_{it}$ and/or $Z_{it}$ through time, Assumption 4 is not restrictive. However, when nonparametric justifications for Assumption 2 (e.g., the exchangeability condition of Altonji and Matzkin (2005)) are involved, Assumption 4 can be restrictive and generally requires that $T$ is larger than $d_{X}$. [^14] One way to justify Assumption 4 when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the $T^{\text{th}}$ order (Altonji and Matzkin (2005)) is to use the symmetry through time in the solution to $\left\{{\bf X}_{i}:V_{it}=v,W_{i}=w\right\}$. To see a concrete example, suppose $T=3$, $X_{it}\in\mathbb{R}^{2}$, $Z_{it}\in\mathbb{R}^{2}$, and $D_{it}\coloneqq\left(X_{it}^{{}^{\prime}},Z_{it}^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathbb{R}^{4}$. If $D_{i}=\left(D_{i1},D_{i2},D_{i3}\right)=\left(a,b,c\right)$ satisfies $V_{it}=v$ for $t=1,2$ and $W_{i}=w$ where all $a,b,c\in\mathbb{R}^{4}$ and the first two coordinates of $a$ and $b$ are linearly independent, by the symmetry of $W_{i}$ in $\left(X_{it},Z_{it}\right)$ through $t$, $\left(b,a,c\right)$ must also satisfy $V_{it}=v$ and $W_{i}=w$ for $t=1,2$. Therefore, given $V_{i1}=v$ and $W_{i}=w$ there are two linearly independent vectors of $\left(a^{\left(1\right)},a^{\left(2\right)}\right)$ and $\left(b^{\left(1\right)},b^{\left(2\right)}\right)$ that $X_{i1}$ can take, satisfying Assumption 4. It is worth pointing out that I only need to use the endogenous coordinates of $X_{it}$ to construct $V_{it}$ and $W_{i}$. Thus, when some coordinates of $X_{it}$ are exogenous in the sense that they are independent of $\left(A_{i},\eta_{it},\varepsilon_{it}\right)$, the unconditional variation in the exogenous coordinates of $X_{it}$ can be exploited to satisfy Assumption 4. See footnote 16 below for more details on this point.
@@
Next, I present another assumption on the residual variation of $X_{it}$ given $V_{it}$ and $W_{i}$ to facilitate the derivative-based argument (3.7).
@@
The support of $X_{it}$ given $V_{it}$ and $W_{i}$ contains some ball of positive radius a.s. with respect to $\left(V_{it},W_{i}\right)$.
Assumption 4 is similar to Assumption 2.2 of Altonji and Matzkin (2005) and is stronger than Assumption 4. When parametric justifications (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2024)) are used for Assumption 2 such that $W_{i}$ contains only mean of $X_{it}$ and/or $Z_{it}$ through time, Assumption 4 is generally not restrictive. However, when nonparametric justifications (e.g., Altonji and Matzkin (2005)) are used for Assumption 2, Assumption 4 typically requires ex-ante information about $W_{i}$ (e.g., the existence of exogenous regressors in (2.1)). This is true because in this case $W_{i}$ includes averages of each endogenous coordinate of $X_{it}$, $Z_{it}$, and their interaction terms up to the $T^{\text{th}}$ order through time, conditioning on which there may not be enough residual variation in $X_{it}$. To address this issue, when nonparametric justifications for Assumption 2 are used, it is advisable to follow the recommendations by Altonji and Matzkin (2005) to ensure continuous residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$. [^15] Specifically, Altonji and Matzkin (2005) suggest to (i) leverage unconditional variation in the exogenous regressors that do not enter $W_{i}$, which I provide details in footnote 16; (ii) restrict the conditional distribution of $A_{i}$ given $W_{i}$ to depend on not all but only a subset of coordinates of $W_{i}$, which is testable by comparing the fit of $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ with $W_{i}$ replaced by a proper subset of it; (iii) assume $f_{\left.A_{i}\right|W_{i}}$ to depend on $W_{i}$ only through a linear combination of the elements of $W_{i}$, i.e., $f_{\left.A_{i}\right|W_{i}}=f_{\left.A_{i}\right|\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}}$ where $W_{i}=\left(W_{i}^{\left(1\right)},...,W_{i}^{\left(d_{W}\right)}\right)^{{}^{\prime}}$ and the $c_{k}$’s are unknown parameters, which implies that $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}\right]$ and the estimation of $c_{k}$’s can proceed by using the method of Ichimura and Lee (1991); (iv) impose a priori restrictions on model (2.1)–(2.2) such that $W_{i}$ enters $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ in a parametric way (e.g., $\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}+\sum_{l=1}^{d_{V}}d_{l}V_{it}^{\left(l\right)}$ when $d_{X}=1$); and, (v) directly impose restrictions on $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$. [^16] To see a concrete example how the unconditional variations in the exogenous regressors help to satisfy Assumption 4, suppose $X_{it}^{\left(1\right)}$ is endogenous while $X_{it}^{\left(-1\right)}$, all other coordinates of $X_{it}$ except $X_{it}^{\left(1\right)}$, are exogenous (independent of $X_{it}^{\left(1\right)}$ and $\beta_{it}$). Suppose the conditional support of $X_{it}^{\left(1\right)}$ given $\left(V_{it},W_{i}\right)$ only contains a non-zero singleton, while the unconditional support of $X_{it}^{\left(-1\right)}$ contains some small ball of positive radius. Denote $b_{1}\left(V_{it},W_{i}\right)=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$. Then, I can identify the first-order moments of $\beta_{it}$ by taking partial derivative with respect to $X_{it}^{\left(-1\right)}$ on both sides of $\mathbb{E}\left[\left.Y_{it}\right|X_{it}^{\left(1\right)},X_{it}^{\left(-1\right)},V_{it},W_{i}\right]=X_{it}^{\left(1\right)}b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)+X_{it}^{\left(-1\right)^{\prime}}b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)$ to identify $b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)$. Finally, identification of $b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)$ can be obtained as $b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.Y_{it}\right|X_{it}^{\left(1\right)},X_{it}^{\left(-1\right)},V_{it},W_{i}\right]-X_{it}^{\left(-1\right)^{\prime}}b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)\right)/X_{it}^{\left(1\right)}.$
@@
Assumption 4 allows the straightforward perturbation-based method (3.7) to identify the APE and LAR without requiring the computation of the inverse of $\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]$. When residual variation is not a concern—such as when $W_{i}$ contains only the mean of $X_{it}$ through time or when exogenous regressors are included in (2.1)—Assumption 4 is preferred for simpler analysis. Accordingly, I impose it for the estimation and inference results in the next section.
I now state my main identification theorem.
@@
If Assumptions 13 and either Assumption 4 or 4 are satisfied, then $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ are identified.
@@ 3.3 Extensions
In my baseline model, I assume that each coordinate of $X_{it}$ depends on only one coordinate of $\eta_{it}$ (as well as $Z_{it}$ and $A_{i}$). Here, I discuss two ways to include vector-valued $\eta_{it}$ into coordinates of $X_{it}$ under additional assumptions motivated by empirical research. To clearly present the idea, I consider the case where $d_{X}=d_{\eta}=2$.
@@
First, timing assumptions on the choice of regressors can be leveraged to allow vector-valued $\eta_{it}$ to enter certain coordinates of $X_{it}$. For instance, in production function applications, choices of later inputs such as capital typically do not depend on random shocks to earlier chosen inputs such as labor and/or material after conditioning on the level of the earlier chose inputs. [^17] Appendix 1 of the 2006 working paper version of Ackerberg, Caves, and Frazer (2015) discusses a similar idea. I thank an anonymous referee for this suggestion. If this timing assumption is imposed, I modify model (2.1)–(2.2) to be
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$
$X_{it}^{\left(1\right)}$ $=g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)}\right),\text{ and }X_{it}^{\left(2\right)}=g^{\left(2\right)}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X_{it}^{\left(1\right)}},Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right).$
Note that now $X_{it}^{\left(2\right)}$ depends on the whole vector of $\eta_{it}$ if $X_{it}^{\left(1\right)}$ is substituted into $g^{\left(2\right)}\left(\cdot\right)$. I propose to use $V_{it}^{\left(2\right)}\coloneqq F_{\left.X_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},Z_{it},W_{i}}\left(\left.X_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},Z_{it},W_{i}\right)$ to control for $\eta_{it}^{\left(2\right)}$ first. Then, I use $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it}^{\left(1\right)}$. The rest of the identification analysis follows as in this paper. [^18] Specifically, for this method to work, I maintain Assumptions 2, 3(a), and 4 (or 4) and change Assumption 1 to be that $g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)}\right)$ and $g^{\left(2\right)}\left(X_{it}^{\left(1\right)},Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$ are strictly monotonic in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively. I also modify Assumption 3(b) to be that $F_{\left.\eta_{it}^{\left(1\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|A_{i},W_{i}\right)$ and $F_{\left.\eta_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},A_{i},W_{i}\right)$ are strictly increasing in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively.
Second, the interaction between different coordinates of $X_{it}$ also provides useful information about certain coordinates of $\eta_{it}$. One way to exploit such information is to assume that while both coordinates of $\eta_{it}$ enter functions that determine each coordinate of $X_{it}$, the ratio of $X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}$ may depend on only one coordinate of $\eta_{it}$, e.g., $\eta_{it}^{\left(2\right)}$ (and $Z_{it}$ and $A_{i}$). For example, in production function applications, the interpretation of $\eta_{it}$ could be that $\eta_{it}^{\left(1\right)}$ is the Hicks neutral productivity and $\eta_{it}^{\left(2\right)}$ is the labor-augmenting technology. Although both material and labor choices depend on both $\eta_{it}^{\left(1\right)}\text{ and }\eta_{it}^{\left(2\right)}$, it is plausible that material per worker depends only on the labor-augmenting technology $\eta_{it}^{\left(2\right)}$. See Proposition 2.1 of Demirer (2022) for more discussions about this assumption. Given this assumption, I can proceed to control for $\eta_{it}^{\left(2\right)}$ first by $V_{it}^{\left(2\right)}\coloneqq F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}\right)$, and then for $\eta_{it}^{\left(1\right)}$ by $V_{it}^{\left(1\right)}\coloneqq F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}\right)$. The rest of the identification analysis follows as in this paper. [^19] Specifically, suppose $X_{it}^{\left(1\right)}=g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)},\eta_{it}^{\left(2\right)}\right)$ and $X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}=g^{\left(2\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$. I maintain Assumptions 2, 3(a), and 4. I modify Assumption 1 to be that $g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)},\eta_{it}^{\left(2\right)}\right)$ and $g^{\left(2\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$ are strictly monotonic in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively. I further change Assumption 3(b) to be $F_{\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}\right)$ and $F_{\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}\right)$ are strictly increasing in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively. Then, I have $V_{it}^{\left(2\right)}\coloneqq$ $F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}\right)$ $=$ $F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},A_{i},W_{i}\right)=F_{\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}\right),$ and $V_{it}^{\left(1\right)}\coloneqq$ $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}\right)$ $=$ $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},A_{i},W_{i}\right)=F_{\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}\right),$ where the second equality holds because by Assumption (2), $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$ and that $V_{it}^{\left(2\right)}$ is a function of $\left(X_{it},Z_{it},W_{i}\right)$ and the last equality holds because $\left(V_{it}^{\left(2\right)},A_{i},W_{i}\right)$ uniquely determines $\eta_{it}^{\left(2\right)}$. Therefore, fixing $\left(X_{it},A_{i},V_{it},W_{i}\right)$ is equivalent to fixing $\left(X_{it},A_{i},\eta_{it},W_{i}\right)$. The rest of the argument follows as in the proof of Theorem 1.

### v3

@@ 3.2 Assumptions
There are at least $d_{X}$ linearly independent points in the support of $X_{it}$ conditional on $(V_{it},W_{i})$ almost surely.
Assumption 4 ensures that $\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]$ in (3.11) is invertible. Thus, the APE and LAR can be identified by (3.11) and the law of iterated expectations.
Assumption 4 is best interpreted as a conditional support (non-degeneracy) requirement: for each $t$, $X_{it}$ must retain sufficient variation conditional on $(V_{it},W_{i})$ to identify the target objects. Requiring residual variation in $X_{it}$ conditional on $V_{it}$ is generally not restrictive (Imbens and Newey (2009)). Essentially, it requires $Z_{it}$ to induce enough variation in $X_{it}$ so that, for each $v\in\text{Supp}(V_{it})$, the set $\left\{X_{it}:V_{it}=v\right\}$ is not a singleton. Therefore, it corresponds to the classical IV relevance condition of $\partial g(z,a,\eta)/\partial z\neq 0$.
The restriction on the support of $X_{it}$ in Assumption 4 mainly comes from conditioning on $W_{i}$, which also concerns Assumption 2. There is an inherent trade-off between Assumptions 2 and 4. To see this, suppose $W_{i}=(\mathbf{X}_{i},\mathbf{Z}_{i})$ which satisfies Assumption 2 trivially. However, because $X_{it}$ is included in $W_{i}$, it makes the support of $X_{it}$ given $W_{i}$ a singleton. Hence, Assumption 4 does not hold unless $X_{it}$ is a scalar and the point in the support of $X_{it}$ given $W_{i}$ is nonzero. More generally, including more elements in $W_{i}$ tends to make Assumption 2 easier to satisfy while making Assumption 4 harder to maintain. Because of this trade-off, I describe next how the conditions used to justify Assumption 2 affect the plausibility of Assumption 4, and give corresponding sufficient conditions.
When justifications for Assumption 2 are used so that $W_{i}$ only includes averages of $X_{it}$ and/or $Z_{it}$ through time (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2025)), Assumption 4 is not restrictive. For instance, with $d_{X}=2$, $T=2$, and $W_{i}=\overline{X}_{i}$, Assumption 4 holds whenever $X_{i1}$ and $X_{i2}$ are not perfectly linearly dependent, so that $\mathrm{Supp}(X_{it}\mid\overline{X}_{i})$ remains non-degenerate. By (1.2) and the relevance condition above, it in turn requires that the serial dependence in $Z_{it}$ cannot be so high that the value of $Z_{it}$ (hence $X_{it}$) for any $t$ is uniquely determined once $W_{i}$ is fixed.
By contrast, when a high-dimensional $W_{i}$ is used to justify Assumption 2 (e.g., via the exchangeability condition of Altonji and Matzkin (2005)), Assumption 4 becomes more restrictive. Intuitively, a high-dimensional $W_{i}$ imposes many constraints on the path $(X_{i1},\ldots,X_{iT})$, so one typically needs longer panel to leave enough feasible configurations and satisfy Assumption 4.
In this case, one possibility is to explore the symmetry through time in the solution to $\left\{{\bf X}_{i}:V_{it}=v,W_{i}=w\right\}$. In particular, when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the $T^{\text{th}}$ order, Assumption 2 is satisfied under a nonparametric exchangeability condition on $f_{\left.\eta_{i}\right|A_{i}}$ by adapting the proof of Altonji and Matzkin (2005) in Proposition 1. Then, since $W_{i}$ is symmetric in $\left(X_{it},Z_{it}\right)$ through time, one may permute the order of $\left(X_{it},Z_{it}\right)$ in time without changing the value of $W_{i}$. This creates enough linearly independent points in the support of $X_{it}$ given $W_{i}$ when $T\geq d_{X}$. When such permutation also does not change the value of $V_{it}$ which is usually satisfied under the relevance condition, Assumption 4 is satisfied. I provide an example in Appendix A.3.
@@
Assumption 4 pertains to the inversion-based argument in (3.11). To facilitate the derivative-based argument in (3.12), I introduce the next assumption, which requires more residual variation of $X_{it}$ conditional on $V_{it}$ and $W_{i}$ than Assumption 4.
@@
The support of $X_{it}$ conditional on $(V_{it},W_{i})$ contains an open ball of positive radius almost surely.
Assumption 4 allows one to perturb $X_{it}$ in $\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]$ when $(V_{it},W_{i})$ is fixed so that conditional moments of $\beta_{it}$ can be recovered by (3.12). It is similar to Assumption 2.2 of Altonji and Matzkin (2005).
The same trade-off between Assumptions 2 and 4 also applies to Assumption 4. When Assumption 2 is justified such that $W_{i}$ is low-dimensional (e.g., $\overline{X}_{i}$ as in Mundlak (1978)), Assumption 4 is generally not restrictive. In essence, $Z_{it}$ must vary enough to generate sufficient residual variation in $X_{it}$ conditional on $(V_{it},W_{i})$.
However, when Assumption 2 is justified using nonparametric arguments (e.g., the exchangeability condition of Altonji and Matzkin (2005)), $d_{W}$ is large and Assumption 4 can be restrictive. To address this issue, I present two sets of solutions in Appendix A.3. First, the unconditional variation of exogenous regressors outside of $W_{i}$ can be leveraged. Second, following the suggestions of Altonji and Matzkin (2005), one may restrict how $W_{i}$ enters $f_{\left.A_{i}\right|W_{i}}$ so that the relevant conditional expectations only concern a subset of $W_{i}$ or a linear index of its components.
@@
Assumption 4 allows the straightforward perturbation-based method (3.12) to identify the APE and LAR without requiring the computation of the inverse of $\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]$. When residual variation is not a concern—such as when $W_{i}$ contains only the mean of $X_{it}$ through time or when exogenous regressors are included in (2.1)—Assumption 4 is preferred for simpler analysis. I compare these two approaches in Remark 2 below.
The next theorem summarizes the main identification result of this paper.
@@
If Assumptions 13 and either Assumption 4 or 4 are satisfied, then the APE $\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}$ and the LAR function $b_{t}(x)=\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right]$ are both identified.
@@ 3.3 Extensions
I present four methods for incorporating vector-valued $\eta_{it}$ into $X_{it}$ under additional assumptions. To illustrate the idea clearly, I assume $d_{X}=d_{\eta}=2$.
@@
First, timing assumptions on the choice of regressors can be exploited. For instance, in production function estimation, choices of later inputs such as capital typically do not depend on random shocks to earlier chosen inputs such as labor after conditioning on the level of the earlier chosen inputs; see Ackerberg, Caves, and Frazer (2015, (Appendix 1 of 2006 Working Paper)) for a similar idea. Under this timing assumption, I modify (2.1)–(2.2) to be:
$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$
$X_{it,1}$ $=g_{1}\left(Z_{it},A_{i},\eta_{it,1}\right),\text{ and }X_{it,2}=g_{2}\left({\color[rgb]{1,0,0}X_{it,1}},Z_{it},A_{i},\eta_{it,2}\right).$ (3.17)
Note that now $X_{it,2}$ depends on the whole vector of $\eta_{it}$ if $X_{it,1}=g_{1}\left(Z_{it},A_{i},\eta_{it,1}\right)$ is substituted into $g_{2}$. I first use $V_{it,2}\coloneqq F_{\left.X_{it,2}\right|X_{it,1},Z_{it},W_{i}}\left(\left.X_{it,2}\right|X_{it,1},Z_{it},W_{i}\right)$ to control for $\eta_{it,2}$ first. Then, I use $F_{\left.X_{it,1}\right|Z_{it},W_{i}}\left(\left.X_{it,1}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it,1}$.
Second, interaction across different coordinates of $X_{it}$ can provide useful information about $\eta_{it}$. One possibility is to assume that, although the full vector $\eta_{it}$ enters each coordinate of $X_{it}$, the ratio $X_{it,1}/X_{it,2}$ depends only on a single component of $\eta_{it}$. For example, in production function estimation, $\eta_{it,1}$ represents the Hicks neutral productivity and $\eta_{it,2}$ denotes the labor-augmenting technology. Although both capital and labor choices depend on vector $\eta_{it}$, it is plausible that capital per worker depends only on the labor-augmenting technology $\eta_{it,2}$. See Proposition 2.1 of Demirer (2025) for more discussions about this assumption. Given this assumption, I can proceed to control for $\eta_{it,2}$ first by
$V_{it,2}\coloneqq F_{\left.X_{it,1}/X_{it,2}\right|Z_{it},W_{i}}\left(\left.X_{it,1}/X_{it,2}\right|Z_{it},W_{i}\right),$ (3.18)
and then for $\eta_{it,1}$ by
$V_{it,1}\coloneqq F_{\left.X_{it,1}\right|Z_{it},V_{it,2},W_{i}}\left(\left.X_{it,1}\right|Z_{it},V_{it,2},W_{i}\right).$ (3.19)
Third, general “global univalence” results—such as Gale and Nikaido (1965) or Palais (1959), Hadamard (1906a, b)—suffice to establish the invertibility required here. In production function estimation, the inverse isotonicity approach in Berry, Gandhi, and Haile (2013) appears promising to obtain global injectivity of $g$ in $\eta_{it}$. In particular, suppose each input choice increases in its own component of $\eta_{it}$, while cross-partials are weakly negative due to substitutability. This happens when, for example, inputs are substitutes in the firm’s optimal input demand; a favorable shock to one input in the form of higher efficiency or lower effective cost leads the firm to substitute toward that input and away from others. Then, the Jacobian matrix of $g$ with respect to $\eta_{it}$, denoted by $J_{g}(\eta_{it})$, has positive diagonal entries and nonpositive off-diagonal entries, i.e., the $Z$-matrix sign pattern required for an $M$-matrix. If, in addition, the own effects are strong enough that the matrix is strictly diagonally dominant, then $J_{g}(\eta_{it})$ is an $M$-matrix. This, in turn, guarantees a unique monotone inverse mapping from observables back to the vector $\eta_{it}$.
Finally, if $\eta_{it}$ enters the structural function $g$ through a scalar index $\eta_{it}^{\prime}\tau$ and $g$ is monotone in that index (Ichimura and Lee (1991)), then the effective unobservable is scalar, satisfying Assumption 1. This aligns with common practice in asset pricing, where a scalar market factor emerges as an index of aggregate shocks and affects stock returns monotonically. A similar structure arises in production function estimation, where scalar productivity summarizes richer underlying states and affects input choices strictly increasingly.